A person’s right eye can see objects clearly only if they are between 25 cm and 75 cm away. (a) What power of contact lens is required so that objects far away are sharp? (b) What will be the near point with the lens in place


1
Expert's answer
2022-03-18T12:01:48-0400

Focal length

"\frac{1}{f}=\frac{1}{u}+\frac{1}{v}"

"\frac{1}{f}=\frac{1}{25}-\frac{1}{75}"

"\frac{1}{u}=\frac{3-1}{75}"

"u=37.5cm"

Power

"P=\frac{1}{F}"

"P=\frac{1}{0.0735}=2.67D"